1. Find the general solution of the differential equation dy/dx = y^2 /1 + x^2


1
Expert's answer
2022-02-02T15:16:25-0500

The given question is; "\displaystyle
\frac{dy}{dx}=\frac{y^2}{1+x^2}"

Now, by method of separation of variables we have;

"\displaystyle
\frac{1}{y^2}\frac{dy}{dx}=\frac{1}{1+x^2}"

integrating both sides with respect to x yields;

"\displaystyle
\int\displaystyle
\frac{1}{y^2}\frac{dy}{dx}\ dx=\int\frac{1}{1+x^2}\ dx\\
\Rightarrow\int y^{-2}\ dy=\int\frac{1}{1+x^2}\ dx\\
\text{But}\int\frac{1}{1+x^2}\ dx=\tan^{-1}(x)\ \text{is standard.}\\
\Rightarrow \frac{y^{-2+1}}{-2+1}=\tan^{-1}(x)+c,\ \text{where c is an arbitrary constant.}\\
\Rightarrow\frac{y^{-1}}{-1}=\tan^{-1}(x)+c\\
\Rightarrow -\frac{1}{y}=\tan^{-1}(x)+c\\
\Rightarrow y=\frac{-1}{\tan^{-1}(x)+c\\}"


Hence, the general solution of the given DE is;

"\displaystyle
y=\frac{-1}{\tan^{-1}(x)+c\\}"